OpenStudy (anonymous):

Use the squared identities to simplify 2cos^2x cos^2x a- 3-4cos(2x)-cos(4x)/4 b- 3+4cos(2x)-cos(4x)/4 c- 3+4cos(2x)+cos(4x)/4 d- 3-4cos(2x)+cos(4x)/4

OpenStudy (anonymous):

=2cos^4x =\[2(\cos ^{2}x)^{2}=2[(1+\cos2x)/2]^{2}\]

OpenStudy (anonymous):

\[1/2(1+2\cos2x+\cos ^{2}2x)=1/2(1+2\cos2x+(1+\cos4x)/2)\]

OpenStudy (anonymous):

\[=1/4(2+4\cos2x+1+\cos4x)=1/4(3+4\cos2x+\cos4x)\]

OpenStudy (anonymous):

Thanks